The logistic growth function
P(x) = 90/1+271e^-0.122x
models the percentage, P(x), of Americans who are x years old with some coronary heard disease.
QUESTION: What percentage of 80-year-olds have some coronary heart disease?

Answer by stanbon(75887)

The logistic growth function
P(x) = 90/1+271e^-0.122x
models the percentage, P(x), of Americans who are x years old with some coronary heard disease.
QUESTION: What percentage of 80-year-olds have some coronary heart disease?
---------------
P(80) = 90/(1+271e^(-0.122*80)
P(80) = 90/1.015640663
P(80) = 88.61%
===============
Cheers,
Stan H.